OpenStudy (anonymous):

1. is -7 + 9 = - 9 + 7 true , false , or open ? ( 1point ) A . true b. false c. open 2. is 4x-3=19 true , false , or open ? ( 1point ) A. true b. false c. open

OpenStudy (igreen):

1. 7 + 9 = - 9 + 7 Add the first side: 16 = -9 + 7 Add the second side: 16 = -2 16 does NOT equal -2, so it is incorrect.

OpenStudy (igreen):

4x - 3 = 19 Ad 3 to both sides: 4x = 22 Divide 4 to both sides: x = 5.5 So the answer is Open, because there can be other values of x, like other fractions.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @diomondsandmakeuo

OpenStudy (anonymous):

so is 1 . false

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Yes, 1 is false, and 2 is open.

OpenStudy (igreen):

@diomondsandmakeuo